OpenStudy (johnnydicamillo):

Find all x-intercepts, minima, and maxima within one period. y=1/2sin((3/4)x)

OpenStudy (johnnydicamillo):

\[y=\frac{ 1 }{ 2 }\sin (\frac{ 3 }{ 4 }x)\]

OpenStudy (mathmate):

The general equation for a sinusoidal function is \(f(x)=a\ sin(b(x-h))+k\) where y=k is the sinusoidal axis, a=1/2, b=3/4 h=0 k=0 Period = \(2\pi /b = 2\pi /(3/4) = 8\pi/3\) max/min = \(k\pm |a| =0\pm 1/2=\pm 1/2 \) 

OpenStudy (johnnydicamillo):

how do I find that question mark? @mathmate

OpenStudy (mathmate):

Hint: The standard sine function is:  Hope that helps.